# Gongxu Luo1,  Yuewen Sun1,2,  Salman Khan1,4,  Peter Spirtes2,  Kun Zhang1,2

### v1

@@ 1 Introduction
Human behavior traits (or behavioral summaries) refer to outwardly observable patterns of conduct inferred from public information such as spoken or written language, facial expressions, and biographical records (Rothe, 2017; Johnson, 1997; DeNeve & Cooper, 1998; Briggs & Cheek, 1986). These traits differ from psychological personality, which concerns internal dispositions typically measured through self-reports or expert evaluation (Cattell et al., 1970; Eysenck & Eysenck, 1975; Myers et al., 1998; Goldberg, 1993). Unlike clinical diagnoses, behavior traits can be inferred ethically and at scale from non-intrusive signals, offering reproducible, population-level insights that complement personality research without medicalizing individuals. Advances in large language models (LLMs) (Achiam et al., 2023; Floridi & Chiriatti, 2020; Jiang et al., 2024; Liu et al., 2024) have further expanded this feasibility. Several studies demonstrate that LLM-based assessments of behavior traits aligned with the Big Five framework can be reliable under carefully designed prompting strategies (Serapio-Garc´ıa et al., 2023; Jiang et al., 2023; Tseng et al., 2024; Zou et al., 2024). These approaches enable large-scale, automated analysis while mitigating some biases inherent in self-reports.
@@
Related Work. Research on human attributes spans two complementary directions: internal personality and external behavior traits. Psychologically internal personality has traditionally been measured with self-report instruments such as the 16PF (Cattell et al., 1970), EPQ (Eysenck & Eysenck, 1975), MBTI (Myers et al., 1998), and the Big Five framework (Goldberg, 1993). In contrast, behavior traits emphasize outwardly observable patterns, inferred from signals such as text, facial expressions, physiology, or digital traces. Several datasets target this perspective, including SALSA for group interactions (Alameda-Pineda et al., 2015), nonsocial-context datasets for daily activities (Dotti et al., 2018), driving and physiological data for trait prediction (Evin et al., 2022), and lifelog corpora capturing multimodal daily behavior (Chung et al., 2022). Digital records such as Facebook Likes have also been shown to predict sensitive traits, including personality dimensions (Kosinski et al., 2013). Other multimodal resources, such as YouTube-Vlogs (Biel & Gatica-Perez, 2012), FI-V2 (Escalante et al., 2020), MuPTA (Ryumina et al., 2023), and MDPE (Cai et al., 2024), combine video, audio, or physiological signals for prediction tasks like impression analysis or deception detection, but they typically lack explicit textual trait descriptions or frameworks for cross-modal interpretation. A comparison table of different datasets is in Tab. A1. Beyond datasets, empirical studies show that observable features in one modality can signal traits in another. For instance, facial structure has been linked to health and aggression cues (Kramer & Ward, 2010; Carré & McCormick, 2008), body images to personality judgments (Naumann et al., 2009), and facial behavior to Big Five traits (Cai & Liu, 2022). Together, these works underscore the promise of behavior trait analysis, but existing resources are limited for systematic cross-modal and causal study. See App. A2 for more details.
@@
[Uncaptioned image] ChatGPT-4o-Latest 4.19 0.03 0.17 0.99 1.00 1.00 1.00 0.96 3.92 0.27 0.17 1.00 1.00 0.99 1.00 0.93
[Uncaptioned image] Gemini-2.5-Pro 23.48 0.06 0.19 0.99 1.00 1.00 1.00 0.96 21.31 0.29 0.22 0.99 1.00 1.00 1.00 0.91
@@
[Uncaptioned image] Llama-4-Maverick 3.73 0.25 0.29 1.00 1.00 0.97 1.00 0.90 3.99 0.30 0.43 1.00 1.00 0.95 1.00 0.87
[Uncaptioned image] Gemini-2.0-Flash-T. 8.83 0.28 0.38 0.99 1.00 1.00 1.00 0.89 8.26 0.48 0.27 0.97 1.00 0.99 1.00 0.87
@@
[Uncaptioned image] Qwen-Plus 9.22 0.24 0.46 1.00 0.99 1.00 1.00 0.88 8.87 0.30 0.48 1.00 0.98 1.00 1.00 0.87
[Uncaptioned image] DeepSeek-V3-0324 14.18 0.47 0.24 0.99 1.00 1.00 1.00 0.88 7.75 0.64 0.20 1.00 1.00 1.00 1.00 0.86
[Uncaptioned image] Gemini-2.0-Flash 2.53 0.43 0.32 1.00 1.00 0.99 1.00 0.87 2.29 0.69 0.18 1.00 1.00 1.00 1.00 0.86
@@
Multimodality. Each record integrates three components: (1) textual behavior-trait descriptions and Big Five scores inferred by LLMs, (2) facial images or embeddings with attribute annotations, and (3) structured biographical metadata. The full feature lists of AthlePersona and CelebPersona are shown in Tab.A2 and Tab.A3, respectively. Dataset distributions for each feature are presented in Fig.LABEL:app_ahtle_distribution and Fig. LABEL:app_celeb_distribution. Terms-of-use compliance across all sports leagues for AthlePersona is summarized in Tab.A4. The complete prompt for inferring behavior traits is provided in Prompt 1.
@@ 3 Analysis Level I: Inferring Statistical Dependence from Structured Data
Figure 3: Independence test (IT) results and distributions of trait scores. (a) and (b) present heatmaps of significant IT results between Big Five behavior traits and other structured features for CelebPersona and AthlePersona, respectively. Each cell reports “x/yx/y,” where $x$ is the number of methods that reject the null hypothesis ($p<0.05$) and $y$ is the total number of applied methods. Lighter shades indicate stronger evidence of dependence. (c) shows the overall distribution of Big Five behavior scores across both datasets. Refer to Tab. LABEL:app_athlete_cit and Tab. LABEL:app_celeb_cit for complete $p$-values.
@@
To examine dependencies between trait scores and other structured features, we apply five different independence test methods: three non-parametric approaches (KCI (Zhang et al., 2012), RCIT (Strobl et al., 2019), HSIC (Gretton et al., 2005)) and two tests designed for discrete variables (Chi-square (Tallarida et al., 1987) and G-square (Tsamardinos et al., 2006)). The detailed descriptions about these methods are summarized in Tab. A6. A dependency is deemed significant if $p<0.05$. Fig.3 shows how many of the five methods found significant dependence, and plots the score distributions. Complete $p$-value results for each independent test method are reported in Tab. LABEL:app_athlete_cit and Tab. LABEL:app_celeb_cit.
@@ 4.2 Identifiability Theory
Assumption i is a moderate condition that ensures the probability distributions are well-defined and computable. Assumption ii informally requires that distinct input distributions correspond to distinct output distributions. In a similar spirit, Assumption iii guarantees that different values of $\mathbf{z}_{m}$ induce different conditional distributions $p(\mathbf{x}_{m,B}\mid\mathbf{z}_{m})$, e.g., heteroskedastic noise. It is important to note that this condition is significantly weaker than monotonicity conditions. Finally, Assumption iv imposes the differentiability of the mapping from $\mathbf{z}_{m},\mathbf{s}]$ to $p_{\mathbf{x}_{m,B}\mid\mathbf{z}_{m},\mathbf{s}}$, which can be explicitly enforced through the use of differentiable models, i.e., variational autoencoders (VAEs).
@@
After identifying the entire latent space and each latent subspace underlying modality observations: $\{[\mathbf{z}_{1},\mathbf{s}]\ldots,[\mathbf{z}_{M},\mathbf{s}]\}$, the shared component $\mathbf{s}$ can be isolated by leveraging the preliminary result that each $\mathbf{s}$ is block-wise identifiable. This enables the application of existing techniques for isolating shared latent spaces, as developed in (Yao et al., 2023; Von Kügelgen et al., 2021).
@@
The core idea is to exploit the rich multi-modal information present in behavior trait datasets to disentangle shared latent variables from modality-specific ones. Shared latent factors, e.g., genetic traits, act as confounders and induce sufficient variability across modality-specific components. This motivates the adoption of nonlinear ICA techniques (Hyvarinen et al., 2019). To achieve identifiability, we impose structural constraints derived from the ground-truth Markov network $\mathcal{M}_{\mathbf{z}_{m}}$ onto the estimated network $\mathcal{M}_{\hat{\mathbf{z}}_{m}}$, leveraging the established connection between conditional independence and vanishing cross-partial derivatives (Lin, 1997): if $z_{m,i}{\perp}z_{m,j}\mid\{\mathbf{s},\mathbf{z}\setminus\{{z_{m,i},z_{m,j}}\}\}$, then $\frac{\partial^{2}\log p(\mathbf{z}_{m})}{\partial z_{m,i}\partial z_{m,j}}=0$. These established conditions shed light on the design of training network.
@@ Acknowledgements
We would also like to acknowledge the support from NSF Award No. 2229881, AI Institute for Societal Decision Making (AI-SDM), the National Institutes of Health (NIH) under Contract R01HL159805, and grants from Quris AI, Florin Court Capital, and MBZUAI-WIS Joint Program.
@@
Human behavior traits have long been central to understanding how individuals reflect on their strengths, limitations, and interpersonal tendencies (Rothe, 2017; Johnson, 1997; DeNeve & Cooper, 1998; Briggs & Cheek, 1986). In contrast to psychological personality, which emphasizes internal dispositions typically assessed through self-reports or expert evaluation, behavior traits focus on outwardly observable patterns inferred from language, facial appearance, physiological states, and digital traces. Traditional self-report instruments, such as Cattell’s 16PF (Cattell et al., 1970), the Eysenck Personality Questionnaire (Eysenck & Eysenck, 1975), and the Myers–Briggs Type Indicator (Myers et al., 1998), provided early frameworks for personality assessment. The Big Five (Goldberg, 1993) has since emerged as the prevailing paradigm, supported by strong empirical evidence and predictive power (Cobb-Clark & Schurer, 2012; Oshio et al., 2018; Komarraju et al., 2011; Roccas et al., 2002; Gerber et al., 2011). However, these methods remain vulnerable to self-report biases and limited scalability.
Computational approaches have sought to infer traits from observable signals rather than introspection. Examples include linguistic cues (Pennebaker et al., 1999), handwriting (Asra & Shubhangi, 2015), speech (Mohammadi et al., 2010), facial expressions (Güçlütürk et al., 2018), and online profiles (Youyou et al., 2015; Celli et al., 2014). A landmark study by (Kosinski et al., 2013) showed that Facebook Likes could predict a wide range of sensitive traits, including personality and demographics, with accuracies comparable to psychometric tests. Beyond digital traces, physiological and behavioral signals have been linked to traits in contexts such as driving (Evin et al., 2022), smart-home daily activities (Dotti et al., 2018), and long-term lifelogging (Chung et al., 2022). These works demonstrate the potential of behavioral data for trait inference, often in settings where traditional questionnaires are impractical.
Several multimodal datasets have been developed to study traits in richer contexts. SALSA (Alameda-Pineda et al., 2015) captures group behavior at social events through multimodal recordings with personality annotations. YouTube-Vlogs (Biel & Gatica-Perez, 2012), FI-V2 (Escalante et al., 2020), MuPTA (Ryumina et al., 2023), MDPE (Cai et al., 2024), and Amigos (Miranda-Correa et al., 2018) integrate video, audio, and physiological data for tasks such as impression analysis or affect recognition. While valuable, these datasets are generally small-scale and lack explicit textual trait descriptions or unified frameworks for cross-modal analysis. Other resources like CelebA (Liu et al., 2015), FFHQ (Karras et al., 2019), and FairFace (Kärkkäinen & Joo, 2021) enable large-scale analysis of facial attributes but do not provide trait or personality annotations.
@@
YouTube-Vlogs (Biel & Gatica-Perez, 2012) Outer Video, Audio, Tabular Personality Prediction
@@
Complementing dataset development, empirical studies have examined how outward features in one modality can signal traits in another. The “kernel of truth” hypothesis (Zebrowitz, 2018) suggests that physical cues may reflect behavioral tendencies, supported by findings linking facial morphology to health (Kramer & Ward, 2010), aggression (Carré & McCormick, 2008), and personality judgments from body images (Naumann et al., 2009). More recently, machine learning methods have shown that Big Five traits can be predicted from facial behavior (Cai & Liu, 2022; Youyou et al., 2015; Kachur et al., 2020). Beyond vision, correlations have also been observed with activity levels (Wilson & Dishman, 2015), sensor data (Dotti et al., 2018), and physiological signals (Gao et al., 2019).
@@
Causal representation learning (CRL) aims to recover high-level causal variables from low-level observations, bridging machine learning and causal inference (Schölkopf et al., 2021; Li et al., 2025; Fan et al., 2024). It generalizes classical causal discovery (Spirtes et al., 2001) by learning structured representations that respect causal semantics. CRL methods with identifiability guarantees typically rely on additional assumptions: (1) Functional constraints on the data-generating process (Xu et al., 2024; Zheng et al., 2022); (2) Interventional or multi-environment data that introduce distributional shifts to expose latent structure (Hyvarinen et al., 2019; Khemakhem et al., 2020); (3) Multimodal or multiview settings, where aligned observations across modalities help identify shared causal factors through sample-level invariance (Yao et al., 2023; Daunhawer et al., 2023; Sturma et al., 2023). Recent studies unify these approaches under general invariance principles, showing that many can be seen as special cases of a broader framework (Ahuja et al., 2023).
Parallel to these theoretical advances, some works focus on practical CRL applications without strict identifiability. Examples include variational methods for biomedical data (Mao et al., 2022), contrastive learning for multimodal causal analysis (Zheng et al., 2024), and causal discovery on neuroimaging datasets (Rawls et al., 2021). In contrast, our work aims to combine both theory and application: we derive formal identifiability conditions under multi-modality multi-measurement settings and integrate these insights into a practical estimation framework for human trait analysis.
@@
Large Language Models (LLMs) such as GPT-4 (OpenAI, 2023), PaLM (Chowdhery et al., 2022), Gemini (Team et al., 2023), DeepSeek (Liu et al., 2024), Qwen (Bai et al., 2023), and Claude (Anthropic, 2023) have demonstrated remarkable reasoning and inference capabilities across a wide range of tasks, including arithmetic (Cobbe et al., 2021), commonsense reasoning (Srivastava et al., 2022), text editing and generation (Tang et al., 2025), code generation (Chen et al., 2021), and scientific QA (Wang et al., 2023). These models perform zero-shot or few-shot reasoning using techniques such as chain-of-thought prompting (Wei et al., 2022), self-consistency (Wang et al., 2023), active prompting (Diao et al., 2023), confidence-based If-or-Else prompting (Li et al., 2024a) and tool-augmented reasoning (Paranjape et al., 2023). Despite their black-box nature, LLMs have shown emergent abilities to perform structured reasoning without explicit supervision, making them powerful general-purpose inference engines.
@@
Recent research has explored the extent to which Large Language Models (LLMs) can infer, simulate, and even express human personality traits. For example, (Peters et al., 2024) demonstrate that models such as GPT-4 can estimate Big Five personality dimensions from user-generated text with moderate accuracy, even in zero-shot settings. Similarly, (Serapio-Garc´ıa et al., 2023) show that LLMs can produce consistent personality profiles when prompted, often aligning with outputs from standardized psychometric assessments. Beyond inference, other studies examine how LLMs naturally exhibit personality-like traits in their responses. (Jiang et al., 2023) introduce methods to control and elicit desired personality traits in language model outputs, while (Wang et al., 2025) analyze the emergent ability of LLMs to emulate distinct personality patterns during generation. Furthermore, recent works such as (Tiuleneva et al., 2024; Rao et al., 2023) utilize LLMs to annotate or assess personality traits from textual data, illustrating their growing role in computational personality research.
@@
NBA Terms of Use $\rightarrow$ 9. NBA STATISTICS (nba, ) “By using such NBA Statistics, you agree that: (i) any use, display, or publication of the NBA Statistics shall include a prominent attribution to NBA.com in connection with such use, display, or publication; (ii) the NBA Statistics may only be used, displayed, or published for legitimate news reporting or private, non-commercial purposes;…” No
@@
NFL Terms and Conditions $\rightarrow$ 1. INTRODUCTION; GENERAL; OWNERSHIP; PROHIBITIONS (nfl, ) “You may use the Services solely for your own individual non-commercial and informational purposes only. Any other use, including for any commercial purposes, is strictly prohibited without our express prior written consent.” No
@@
MLB Terms of Use Agreement (mlb, ) “… you must not reproduce, prepare derivative works based upon, distribute, perform or display the MLB Digital Properties without first obtaining the written permission of MLB or otherwise as expressly set forth in the terms and conditions of the MLB Digital Properties. The MLB Digital Properties must not be used in any unauthorized manner.” Yes
@@
NHL Terms of Service $\rightarrow$ 7. Intellectual Property (nhl, ) “You may access, use, and display the Services, but only for non-commercial, informational, personal use, without modification or alteration in any way, and only so long as you comply with these Terms.” No
@@
Premier League Terms of Use $\rightarrow$ Intellectual Property Rights (pre, ) “You may download and print material from the Website or App as is reasonable for your own private and personal use. You may also forward such material from the Website or App to other people for their private and personal use provided you credit us as its source and add the Website address.” No
@@
La Liga Legal Notice and Conditions of Use $\rightarrow$ 3. Use of the Website(lal, ) “… The User undertakes to refrain from (a) using the Contents in a manner… (b) reproduce or copy, distribute, allow public access through any form of public communication, adapt, transform or modify the Contents, unless authorised by the owner of the corresponding rights or it is legally permitted…” Yes
@@
Serie A General Terms and Conditions of the License Agreement $\rightarrow$ 2. Right limitations $\rightarrow$ 2.2 (ii) Official Data(leg, ) “Except in the case of a separate written agreement between Lega Serie A and the Licensee establishing otherwise, the Licensee may only exploit the data related to the Competitions, the Matches, the Clubs and the players in the context …” Yes

### v2

@@ 1 Introduction
Human behavior traits (or behavioral summaries) refer to outwardly observable patterns of conduct inferred from public information such as spoken or written language, facial expressions, and biographical records (Rothe, 2017; Johnson, 1997; DeNeve and Cooper, 1998; Briggs and Cheek, 1986). These traits differ from psychological personality, which concerns internal dispositions typically measured through self-reports or expert evaluation (Cattell et al., 1970; Eysenck and Eysenck, 1975; Myers et al., 1998; Goldberg, 1993). Unlike clinical diagnoses, behavior traits can be inferred ethically and at scale from non-intrusive signals, offering reproducible, population-level insights that complement personality research without medicalizing individuals. Advances in large language models (LLMs) (Achiam et al., 2023; Floridi and Chiriatti, 2020; Jiang et al., 2024; Liu et al., 2024) have further expanded this feasibility. Several studies demonstrate that LLM-based assessments of behavior traits aligned with the Big Five framework can be reliable under carefully designed prompting strategies (Serapio-García et al., 2023; Jiang et al., 2023; Tseng et al., 2024; Zou et al., 2024). These approaches enable large-scale, automated analysis while mitigating some biases inherent in self-reports.
@@
Related Work. Research on human attributes spans two complementary directions: internal personality and external behavior traits. Psychologically internal personality has traditionally been measured with self-report instruments such as the 16PF (Cattell et al., 1970), EPQ (Eysenck and Eysenck, 1975), MBTI (Myers et al., 1998), and the Big Five framework (Goldberg, 1993). In contrast, behavior traits emphasize outwardly observable patterns, inferred from signals such as text, facial expressions, physiology, or digital traces. Several datasets target this perspective, including SALSA for group interactions (Alameda-Pineda et al., 2015), nonsocial-context datasets for daily activities (Dotti et al., 2018), driving and physiological data for trait prediction (Evin et al., 2022), and lifelog corpora capturing multimodal daily behavior (Chung et al., 2022). Digital records such as Facebook Likes have also been shown to predict sensitive traits, including personality dimensions (Kosinski et al., 2013). Other multimodal resources, such as YouTube-Vlogs (Biel and Gatica-Perez, 2012), FI-V2 (Escalante et al., 2020), MuPTA (Ryumina et al., 2023), and MDPE (Cai et al., 2024), combine video, audio, or physiological signals for prediction tasks like impression analysis or deception detection, but they typically lack explicit textual trait descriptions or frameworks for cross-modal interpretation. A comparison table of different datasets is in Tab. A1. Beyond datasets, empirical studies show that observable features in one modality can signal traits in another. For instance, facial structure has been linked to health and aggression cues (Kramer and Ward, 2010; Carré and McCormick, 2008), body images to personality judgments (Naumann et al., 2009), and facial behavior to Big Five traits (Cai and Liu, 2022). Together, these works underscore the promise of behavior trait analysis, but existing resources are limited for systematic cross-modal and causal study. See App. A2 for more details.
@@
[Uncaptioned image] ChatGPT-4o 4.19 0.03 0.17 0.99 1.00 1.00 1.00 0.96 3.92 0.27 0.17 1.00 1.00 0.99 1.00 0.93
[Uncaptioned image] Gemini2.5-Pro 23.48 0.06 0.19 0.99 1.00 1.00 1.00 0.96 21.31 0.29 0.22 0.99 1.00 1.00 1.00 0.91
@@
[Uncaptioned image] Llama-4 3.73 0.25 0.29 1.00 1.00 0.97 1.00 0.90 3.99 0.30 0.43 1.00 1.00 0.95 1.00 0.87
[Uncaptioned image] Gemini2.0-FT 8.83 0.28 0.38 0.99 1.00 1.00 1.00 0.89 8.26 0.48 0.27 0.97 1.00 0.99 1.00 0.87
@@
[Uncaptioned image] QwQ-32B 9.22 0.24 0.46 1.00 0.99 1.00 1.00 0.88 8.87 0.30 0.48 1.00 0.98 1.00 1.00 0.87
[Uncaptioned image] DeepSeek-V3 14.18 0.47 0.24 0.99 1.00 1.00 1.00 0.88 7.75 0.64 0.20 1.00 1.00 1.00 1.00 0.86
[Uncaptioned image] Gemini2.0-F 2.53 0.43 0.32 1.00 1.00 0.99 1.00 0.87 2.29 0.69 0.18 1.00 1.00 1.00 1.00 0.86
@@
Multimodality. Each record integrates three components: (1) textual behavior-trait descriptions and Big Five scores inferred by LLMs, (2) facial images or embeddings with attribute annotations, and (3) structured biographical metadata. The full feature lists of AthlePersona and CelebPersona are shown in Tab.A2 and Tab.A3, respectively. Dataset distributions for each feature are presented in Fig.7(a) and Fig. 7(b). Terms-of-use compliance across all sports leagues for AthlePersona is summarized in Tab.A4. The complete prompt for inferring behavior traits is provided in Prompt 1.
@@ 3 Analysis Level I: Inferring Statistical Dependence from Structured Data
Figure 3: Independence test (IT) results and distributions of trait scores. (a) and (b) present heatmaps of significant IT results between Big Five behavior traits and other structured features for CelebPersona and AthlePersona, respectively. Each cell reports “x/yx/y,” where $x$ is the number of methods that reject the null hypothesis ($p<0.05$) and $y$ is the total number of applied methods. Lighter shades indicate stronger evidence of dependence. (c) shows the overall distribution of Big Five behavior scores across both datasets. Refer to Tab. 13(a) and Tab. 13(b) for complete $p$-values.
@@
To examine dependencies between trait scores and other structured features, we apply five different independence test methods: three non-parametric approaches (KCI (Zhang et al., 2012), RCIT (Strobl et al., 2019), HSIC (Gretton et al., 2005)) and two tests designed for discrete variables (Chi-square (Tallarida et al., 1987) and G-square (Tsamardinos et al., 2006)). The detailed descriptions about these methods are summarized in Tab. A6. A dependency is deemed significant if $p<0.05$. Fig.3 shows how many of the five methods found significant dependence, and plots the score distributions. Complete $p$-value results for each independent test method are reported in Tab. 13(a) and Tab. 13(b).
@@ 4.2 Identifiability Theory
Assumption i is a moderate condition that ensures the probability distributions are well-defined and computable. Assumption ii informally requires that distinct input distributions correspond to distinct output distributions. In a similar spirit, Assumption iii guarantees that different values of $\mathbf{z}_{m}$ induce different conditional distributions $p(\mathbf{x}_{m,B}\mid\mathbf{z}_{m})$, e.g., heteroskedastic noise. Notably, this condition is significantly weaker than monotonicity. Finally, Assumption iv imposes the differentiability of the mapping from $\mathbf{z}_{m},\mathbf{s}]$ to $p_{\mathbf{x}_{m,B}\mid\mathbf{z}_{m},\mathbf{s}}$, which can be explicitly enforced through the use of differentiable models, i.e., variational autoencoders (VAEs).
@@
After identifying the entire latent space and each latent subspace underlying modality observations: $\{[\mathbf{z}_{1},\mathbf{s}]\ldots,[\mathbf{z}_{M},\mathbf{s}]\}$, the shared component $\mathbf{s}$ can be isolated by leveraging the preliminary result that each $\mathbf{s}$ is block-wise identifiable. This enables the application of existing techniques for isolating shared latent spaces, e.g., as developed in (Yao et al., 2023; Von Kügelgen et al., 2021).
@@
The core idea is to exploit the rich multi-modal information present in behavior trait datasets to disentangle shared latent variables from modality-specific ones. Shared latent factors, e.g., genetic traits, act as confounders and induce sufficient variability across modality-specific components. This motivates the adoption of nonlinear ICA (Hyvarinen et al., 2019). To achieve identifiability, we impose structural constraints derived from the ground-truth Markov network $\mathcal{M}_{\mathbf{z}_{m}}$ onto the estimated network $\mathcal{M}_{\hat{\mathbf{z}}_{m}}$, leveraging the connection between conditional independence and vanishing cross-partial derivatives (Lin, 1997): if $z_{m,i}{\perp}z_{m,j}\mid\{\mathbf{s},\mathbf{z}\setminus\{{z_{m,i},z_{m,j}}\}\}$, then $\frac{\partial^{2}\log p(\mathbf{z}_{m})}{\partial z_{m,i}\partial z_{m,j}}=0$. These established conditions shed light on the design of training network below.
@@ Acknowledgements
We would also like to acknowledge the support from NSF Award No. 2229881, AI Institute for Societal Decision Making (AI-SDM), the National Institutes of Health (NIH) under Contract R01HL159805, and grants from Quris AI, Florin Court Capital, MBZUAI-WIS Joint Program, and the Al Deira Causal Education project.
@@
Human behavior traits have long been central to understanding how individuals reflect on their strengths, limitations, and interpersonal tendencies (Rothe, 2017; Johnson, 1997; DeNeve and Cooper, 1998; Briggs and Cheek, 1986). In contrast to psychological personality, which emphasizes internal dispositions typically assessed through self-reports or expert evaluation, behavior traits focus on outwardly observable patterns inferred from language, facial appearance, physiological states, and digital traces. Traditional self-report instruments, such as Cattell’s 16PF (Cattell et al., 1970), the Eysenck Personality Questionnaire (Eysenck and Eysenck, 1975), and the Myers–Briggs Type Indicator (Myers et al., 1998), provided early frameworks for personality assessment. The Big Five (Goldberg, 1993) has since emerged as the prevailing paradigm, supported by strong empirical evidence and predictive power (Cobb-Clark and Schurer, 2012; Oshio et al., 2018; Komarraju et al., 2011; Roccas et al., 2002; Gerber et al., 2011). However, these methods remain vulnerable to self-report biases and limited scalability.
Computational approaches have sought to infer traits from observable signals rather than introspection. Examples include linguistic cues (Pennebaker et al., 1999), handwriting (Asra and Shubhangi, 2015), speech (Mohammadi et al., 2010), facial expressions (Güçlütürk et al., 2018), and online profiles (Youyou et al., 2015; Celli et al., 2014). A landmark study by (Kosinski et al., 2013) showed that Facebook Likes could predict a wide range of sensitive traits, including personality and demographics, with accuracies comparable to psychometric tests. Beyond digital traces, physiological and behavioral signals have been linked to traits in contexts such as driving (Evin et al., 2022), smart-home daily activities (Dotti et al., 2018), and long-term lifelogging (Chung et al., 2022). These works demonstrate the potential of behavioral data for trait inference, often in settings where traditional questionnaires are impractical.
Several multimodal datasets have been developed to study traits in richer contexts. SALSA (Alameda-Pineda et al., 2015) captures group behavior at social events through multimodal recordings with personality annotations. YouTube-Vlogs (Biel and Gatica-Perez, 2012), FI-V2 (Escalante et al., 2020), MuPTA (Ryumina et al., 2023), MDPE (Cai et al., 2024), and Amigos (Miranda-Correa et al., 2018) integrate video, audio, and physiological data for tasks such as impression analysis or affect recognition. While valuable, these datasets are generally small-scale and lack explicit textual trait descriptions or unified frameworks for cross-modal analysis. Other resources like CelebA (Liu et al., 2015), FFHQ (Karras et al., 2019), and FairFace (Kärkkäinen and Joo, 2021) enable large-scale analysis of facial attributes but do not provide trait or personality annotations.
@@
YouTube-Vlogs (Biel and Gatica-Perez, 2012) Outer Video, Audio, Tabular Personality Prediction
@@
Complementing dataset development, empirical studies have examined how outward features in one modality can signal traits in another. The “kernel of truth” hypothesis (Zebrowitz, 2018) suggests that physical cues may reflect behavioral tendencies, supported by findings linking facial morphology to health (Kramer and Ward, 2010), aggression (Carré and McCormick, 2008), and personality judgments from body images (Naumann et al., 2009). More recently, machine learning methods have shown that Big Five traits can be predicted from facial behavior (Cai and Liu, 2022; Youyou et al., 2015; Kachur et al., 2020). Beyond vision, correlations have also been observed with activity levels (Wilson and Dishman, 2015), sensor data (Dotti et al., 2018), and physiological signals (Gao et al., 2019).
@@
Causal representation learning (CRL) aims to recover high-level causal variables from low-level observations, bridging machine learning and causal inference (Schölkopf et al., 2021; Li et al., 2025a; Fan et al., 2024). It generalizes classical causal discovery (Spirtes et al., 2001) by learning structured representations that respect causal semantics. CRL methods with identifiability guarantees typically rely on additional assumptions: (1) Functional constraints on the data-generating process (Xu et al., 2024; Zheng et al., 2022); (2) Interventional or multi-environment data that introduce distributional shifts to expose latent structure (Hyvarinen et al., 2019; Khemakhem et al., 2020); (3) Multimodal or multiview settings, where aligned observations across modalities help identify shared causal factors through sample-level invariance (Yao et al., 2023; Daunhawer et al., 2023; Sturma et al., 2023). Recent studies unify these approaches under general invariance principles, showing that many can be seen as special cases of a broader framework (Ahuja et al., 2023).
Parallel to these theoretical advances, some works focus on practical CRL applications without strict identifiability. Examples include variational methods for biomedical data (Mao et al., 2022), contrastive learning for multimodal causal analysis (Zheng et al., 2024), learning causal insights from OpenReview systems (Li et al., ), and causal discovery on neuroimaging datasets (Rawls et al., 2021). In contrast, our work aims to combine both theory and application: we derive formal identifiability conditions under multi-modality multi-measurement settings and integrate these insights into a practical estimation framework for human trait analysis.
@@
Large Language Models (LLMs) such as GPT-4 (OpenAI, 2023), PaLM (Chowdhery et al., 2022), Gemini (Team et al., 2023), DeepSeek (Liu et al., 2024), Qwen (Bai et al., 2023), and Claude (Anthropic, 2023) have demonstrated remarkable reasoning and inference capabilities across a wide range of tasks, including arithmetic (Cobbe et al., 2021), commonsense reasoning (Srivastava and others, 2022), text editing and generation (Tang et al., 2025), code generation (Chen et al., 2021), and scientific QA (Wang et al., 2023). These models perform zero-shot or few-shot reasoning using techniques such as chain-of-thought prompting (Wei et al., 2022), self-consistency (Wang et al., 2023), active prompting (Diao et al., 2023), confidence-based If-or-Else prompting (Li et al., 2024a), reflective reasoning (Du et al., 2025) and tool-augmented reasoning (Paranjape and others, 2023). Despite their black-box nature, LLMs have shown emergent abilities to perform structured reasoning without explicit supervision, making them powerful general-purpose inference engines.
@@
Recent research has explored the extent to which Large Language Models (LLMs) can infer, simulate, and even express human personality traits. For example, (Peters et al., 2024) demonstrate that models such as GPT-4 can estimate Big Five personality dimensions from user-generated text with moderate accuracy, even in zero-shot settings. Similarly, (Serapio-García et al., 2023) show that LLMs can produce consistent personality profiles when prompted, often aligning with outputs from standardized psychometric assessments. Beyond inference, other studies examine how LLMs naturally exhibit personality-like traits in their responses. (Jiang et al., 2023) introduce methods to control and elicit desired personality traits in language model outputs, while (Wang et al., 2025) analyze the emergent ability of LLMs to emulate distinct personality patterns during generation. Furthermore, recent works such as (Tiuleneva et al., 2024; Rao et al., 2023) utilize LLMs to annotate or assess personality traits from textual data, illustrating their growing role in computational personality research.
@@
NBA Terms of Use $\rightarrow$ 9. NBA STATISTICS (88) “By using such NBA Statistics, you agree that: (i) any use, display, or publication of the NBA Statistics shall include a prominent attribution to NBA.com in connection with such use, display, or publication; (ii) the NBA Statistics may only be used, displayed, or published for legitimate news reporting or private, non-commercial purposes;…” No
@@
NFL Terms and Conditions $\rightarrow$ 1. INTRODUCTION; GENERAL; OWNERSHIP; PROHIBITIONS (89) “You may use the Services solely for your own individual non-commercial and informational purposes only. Any other use, including for any commercial purposes, is strictly prohibited without our express prior written consent.” No
@@
MLB Terms of Use Agreement (84) “… you must not reproduce, prepare derivative works based upon, distribute, perform or display the MLB Digital Properties without first obtaining the written permission of MLB or otherwise as expressly set forth in the terms and conditions of the MLB Digital Properties. The MLB Digital Properties must not be used in any unauthorized manner.” Yes
@@
NHL Terms of Service $\rightarrow$ 7. Intellectual Property (91) “You may access, use, and display the Services, but only for non-commercial, informational, personal use, without modification or alteration in any way, and only so long as you comply with these Terms.” No
@@
Premier League Terms of Use $\rightarrow$ Intellectual Property Rights (102) “You may download and print material from the Website or App as is reasonable for your own private and personal use. You may also forward such material from the Website or App to other people for their private and personal use provided you credit us as its source and add the Website address.” No
@@
La Liga Legal Notice and Conditions of Use $\rightarrow$ 3. Use of the Website(62) “… The User undertakes to refrain from (a) using the Contents in a manner… (b) reproduce or copy, distribute, allow public access through any form of public communication, adapt, transform or modify the Contents, unless authorised by the owner of the corresponding rights or it is legally permitted…” Yes
@@
Serie A General Terms and Conditions of the License Agreement $\rightarrow$ 2. Right limitations $\rightarrow$ 2.2 (ii) Official Data(64) “Except in the case of a separate written agreement between Lega Serie A and the Licensee establishing otherwise, the Licensee may only exploit the data related to the Competitions, the Matches, the Clubs and the players in the context …” Yes